OpenStudy (anonymous):

Express the following function, F(x) as a composition of two functions f and g. F(x)= x^2/(x^2 + 4)

OpenStudy (ranga):

There are many solutions for f and g. Here is one solution: \[F(x) = \frac{ x ^{2} }{ (x ^{2} + 4)}\] \[g(x) = x ^{2}\] and \[f(x) = \frac{ x }{(x + 4) }\] So F(x) = (f o g)(x)

OpenStudy (anonymous):

Would I write f(x) (f o g)(x) as my answer or include the x^2 and x/(x+4) too so it shows my reasoning?

OpenStudy (ranga):

You can write the answer as: g(x) = x^2 f(x) = x / (x + 4) F(x) = (f o g)(x) = x^2 / (x^2 + 4)

OpenStudy (ranga):

Don't forget the original function is upper case F and the two functions are lower case f and g as the problem states.

OpenStudy (anonymous):

Right. Ok, seriously thank you. Hopefully I do well on this upcoming test

OpenStudy (ranga):

Best Wishes!